# Exploring the Frontiers of Softmax: Provable Optimization, Applications in Diffusion Model, and Beyond

### v1

@@ 5 Proof Sketch
Then with probability at least $1-\delta$, we have
$\|H(W)-H(\widetilde{W})\|_{F}\leq Rnd\exp(10B).$
Please see Appendix B.2 for the proof of Lemma 5.1. We can see that the kernel matrix has a small perturbation when the weights $w$ perturb. Note that in Lemma 4.2 [94], they have $\|H(W)-H(\widetilde{W})\|_{F}\leq 2Rn$ for the ReLU activation function and in Lemma 6.7 [53], they have $\|H(W)-H(\widetilde{W})\|_{F}\leq 3Rn^{1+o(1)}$ for the $\exp$ activation function. When we consider the 1-dimension linear regression task, we have $\|H(W)-H(\widetilde{W})\|_{F}\leq Rn^{1+o(1)}$, which is almost the same as the other two cases.
@@
Lemma 5.3 means that, at a fixed timestamp $\tau$, if the weights $w(\tau)$ is close to its initialization, the loss is decreasing and the gradient is also small, then we can conclude at timestamp $\tau+1$, these conditions still hold as local convexity proved by Lemma 5.1. Thus, after checking the initial condition, we can conclude Theorem 4.2.
6 Application in Diffusion
@@ 6.1 Preliminary of Diffusion
$x(t)=e^{-\int_{0}^{t}\frac{1}{2}g(s)\mathrm{d}s}x(0)+e^{-\int_{0}^{t}\frac{1}{2}g(s)\mathrm{d}s}\int_{0}^{t}e^{\int_{0}^{s}\frac{1}{2}g(u)\mathrm{d}u}\sqrt{g(s)}\mathrm{d}B_{s}.$
Backward Process. We denote $y(t)=x(T-t)$ to reverse the forward process in time [65, 47, 20] that transforms noise into samples from the target distribution. We have a backward process associated to Eq. (3) as:
$\mathrm{d}y(t)=(\frac{1}{2}g(T-t)y(t)+g(T-t)\nabla\log p_{T-t}(y(t)))\mathrm{d}t+\sqrt{g(T-t)}\mathrm{d}\bar{B}_{t},\quad y(0)\sim q_{0}.$ (4)
where $(\bar{B}_{t})_{t\geq 0}$ is another $d$-dim Brownian motion/noise. Following the literature, we call $\nabla\log p_{t}(\cdot)$ as “score function” [129]. We have $q_{0}$ is the initial distribution of the backward process and the score function $\nabla\log p_{t}(\cdot)$ as the gradient of log density of $x(t)$.
However, In practice, Eq.(4) cannot be directly used as both the score function and the distribution $p_{T}$ are unknown. To solve the problem, we (1) randomly select a noise distribution as the initial distribution of the backward process $p_{T}$; (2) replace the ground-truth score function $\nabla\log p_{t}(x(t))$ by an estimator $s_{\theta}(x(t),t)$. The parameterized estimator $s_{\theta}$ is learned by a neural network such as U-Net [60, 112] and Transformer [107]. Thus, we obtain a practically implementable approximation of the backward SDE:
$\mathrm{~{}d}y(t)=(\frac{1}{2}g(T-t)y(t)+g(T-t)s_{\theta}(y(t),t))\mathrm{d}t+\sqrt{g(T-t)}\mathrm{d}\bar{B}_{t},\quad y(0)\sim\mathcal{N}(0,I_{d}),$
@@
Score Matching. When estimate the score function, usually we use $L_{2}$ loss between the estimated and actual score:
@@
$\min_{\theta}\frac{1}{T-T_{0}}\int_{T_{0}}^{T}\lambda(t)\mathbb{E}[\|s_{\theta}(x(t),t)-\nabla\log p_{t\mid 0}(x(t)\mid x(0))\|_{2}^{2}]\mathrm{d}t.$ (6)
In practice, the estimator of the score function is parameterized by a neural network and we have the following sampling procedure for any $i\in[n]$,
@@
and we get the training dataset $\{x(0)_{i},(t_{i},x(t_{i})_{i})\}_{i=1}^{n}$, where $x(0)_{i}\in\mathbb{R}^{d}$ and $(t_{i},x(t_{i})_{i})\in\mathbb{R}^{d+1}$. We denote $x(0)$ as the noisy label and $\operatorname*{{\mathbb{E}}}[x(0)|x(t)]$ as the true label. For simplicity, we denote $x(0)_{i}$ as $y_{i}\in\mathbb{R}^{d}$ and $(t_{i},x(t_{i})_{i})$ as $x_{i}\in\mathbb{R}^{d+1}$ and the training dataset as ${\cal D}_{n}=\{(x_{i},y_{i})\}_{i=1}^{n}$. Here, $y$ denotes the image from a dataset and $x$ denotes the noised image with its diffusion process time $t$.
Neural Network Parameterization. Recall that we consider a two-layer network with softmax activation function as the diffusion model in Eq. (1), satisfying $\forall\ell\in[d]$, $F(W,x,a)_{\ell}={{}{}{}m}\langle a_{\ell},\exp(W^{\top}x)\rangle\cdot\langle\exp(W^{\top}x),{\bf 1}_{m}\rangle^{-1}$. Note that, we do not train the top-layer weights $a$, so we can denote it as $F_{nn}(W,x)$.
Then, similar as [60, 66], our loss function Eq. (6) can be rewrite as
@@ 6.2 Main Result of Diffusion
The function $g$ is almost everywhere continuous and bounded on $[0,\infty)$.
@@
$\Gamma_{\delta}:=$ $\left(\frac{2d^{2}A(R_{\mathcal{H}})}{\lambda}\log^{3/2}(\frac{e(dn)^{3/2}A(R_{\mathcal{H}})}{\lambda})+\frac{1}{\sqrt{n}}\right)^{2}+\frac{d^{2}A^{2}(R_{\mathcal{H}})}{\lambda^{2}}(\log(1/\delta)+\log(\log n)).$
@@
Suppose Assumptions 6.16.26.36.4 hold and we set $m=\Omega(\lambda^{-2}n^{3}d^{3}\exp(18B)\log^{2}(nd/\delta))$ and $\eta=0.1\lambda/(mn^{2}d^{2}\exp(16B))$. Moreover, suppose $\widehat{T}$ satisfies Assumption G.3 with corresponding $\epsilon(n,\widehat{T})$. Then for large enough $R_{\mathcal{H}}$, with probability at least $1-\delta$, it holds that
Please refer to Appendix G.1 for the complete proof. Here we provide a proof sketch.
@@
7 Discussion and Future Work
Self-attention Learning. The self-attention can be written as
$F(W^{K}X,W^{Q}X,W^{V}X)\in\mathbb{R}^{d\times n^{\prime}},$ (7)
Feature Learning. Recently, there is a line of work showing that feature learning may be beyond NTK on sample complexity or time complexity, e.g., [9, 149, 63, 10, 43, 18, 155, 71, 85, 52, 115, 92, 90, 40, 133, 134] and many more. It is worth studying the feature learning ability of two-layer softmax NN to figure out what feature pattern the softmax prefers to learn and how it happens. We leave this valuable direction as a future work.
@@ 8 Conclusion
This paper provides a theoretical analysis of the optimization and generalization properties of two-layer neural networks with softmax activation function. We apply our results in learning score estimation functions in diffusion models with noisy labels to verify our analysis effectiveness. Our findings contribute to a deeper understanding of the power of softmax neural networks and their potential to self-attention, advance LLMs, and generative modeling.
@@ Acknowledgement
Roadmap. In Section A, we introduce some definitions that will be used in the proof. In Section B, we provide the basic concentration. In Section C, we provide the proof of our inductions. In Section D, we establish a bound for the weight of induction Part 1. In Section E, we establish a bound for the loss of induction Part 2. In Section F, we introduce the NTK regression. In Section G, we introduce the diffusion.
@@ C.1 Main Result
Then, after $\widehat{T}$ iterations, we have
@@ E.3 Bounding $C_{0}$
$=$ $~{}{{}{}{}m}\sum_{r=1}^{m}a_{\ell,r}\alpha_{i}(\tau+1)^{-1}\alpha_{i}(\tau)^{-1}\sum_{r_{2}=1}^{m}-\eta\langle\Delta w_{r_{2}}(\tau),x_{i}\rangle\exp(w_{r_{2}}(\tau)x_{i})\exp(\langle w_{r}(\tau+1),x_{i}\rangle)$
@@ Appendix F NTK Regression
Recall that, by Definition 3.6, we have the $H^{*}=H(W(0))$. Recall $[H^{*}]_{i,j}\in\mathbb{R}^{d\times d}$ is the kernel between $x_{i}$ and $x_{j}$. Similarly, $\forall\ell_{1},\ell_{2}\in[d]$, for test data, we can define the NTK induced feature map as
@@ F.1 Equivalence between Trained Net and Kernel Regression
where we have used the fact that the initial prediction is 00 as $\beta(0)={\bf 0}_{d}$. Similarly for $F_{nn}(x_{te})_{\ell}$. Let $F_{ntk2,i}(\tau)=F_{ntk2}(\beta(\tau),x_{i})$ and $F_{ntk2}(\tau)\in\mathbb{R}^{d\times n}$. Similarly, for the NN predictor $F_{nn}$. Now we take a closer look at the time derivative:
@@
Since $H^{*}$ is positive semidefinite, $-H^{*}\operatorname{vec}(F_{nn}(\tau)-F_{ntk2}(\tau))$ term only makes $\|F_{nn}(\tau)-F_{ntk2}(\tau)\|_{F}$ smaller. Therefore, we have
@@ Appendix G Diffusion
[…]
@@ G.1 Main Result of Diffusion
Suppose Assumptions 6.16.26.36.4 hold and we set $m=\Omega(\lambda^{-2}n^{3}d^{3}\exp(18B)\log^{2}(nd/\delta))$ and $\eta=0.1\lambda/(mn^{2}d^{2}\exp(16B))$. Moreover, suppose $\widehat{T}$ satisfies Assumption G.3 with corresponding $\epsilon(n,\widehat{T})$. Then for large enough $R_{\mathcal{H}}$, with probability at least $1-\delta$, it holds that
@@
Thus, we finish the proof by Claim G.1, where coupling is from above, label mismatch is from Theorem G.5, early stopping is from Assumption G.3 and approximation is from Theorem G.4. ∎
@@ G.2 Tools From Previous Works
We have the following assumption and statements from previous works [66].
Assumption G.3 (Assumption 3.11 in [66]).
Fix any $F_{\mathcal{H}}\in\mathcal{H}$ with $\|F_{\mathcal{H}}\|_{\mathcal{H}}^{2}\leq R_{\mathcal{H}}$ and assume labels are generated as $\widetilde{y}_{j}=F_{\mathcal{H}}(x_{j})+\epsilon_{j}$. Suppose $\widetilde{F}_{ntk}(\gamma(\widehat{T}),\cdot)$ is obtained by GD-trained kernel regression with the number of iterations $\widehat{T}$. We assume there exists $\epsilon$ such that
$\frac{1}{T}\int_{0}^{T}\mathbb{E}[\widetilde{F}_{ntk}(\gamma(\widehat{T}),(t,x(t)))-F_{\mathcal{H}}(t,x(t))\|_{2}^{2}]\mathrm{d}t\leq\epsilon(n,\widehat{T}),$
and $\epsilon(n,\widehat{T})\rightarrow 0$ as $n\rightarrow\infty.$
@@
$\frac{1}{T}\int_{0}^{T}\mathbb{E}[\|F_{ntk}(\gamma(\tau),(t,x(t)))-\widetilde{F}_{ntk}(\gamma(\tau),(t,x(t)))\|_{2}^{2}]\mathrm{d}t\leq dA(R_{\mathcal{H}})+C_{0}(\sqrt{dA(R_{\mathcal{H}})\Gamma_{\delta}}+\Gamma_{\delta})$

### v2

@@ 5 Technical Overview
Then, with probability at least $1-\delta$, we have $\|H(W)-H(\widetilde{W})\|_{F}\leq Rnd\exp(10B).$
Please see Appendix B.2 for the proof of Lemma 5.1. We can see that the kernel matrix has a small perturbation when the weights $w$ perturb. Note that in Lemma 4.2 of [97], they have $\|H(W)-H(\widetilde{W})\|_{F}\leq 2Rn$ for the ReLU activation function and in Lemma 6.7 of [48], they have $\|H(W)-H(\widetilde{W})\|_{F}\leq 3Rn^{1+o(1)}$ for the $\exp$ activation function. When we consider the 1-dimension linear regression task, we have $\|H(W)-H(\widetilde{W})\|_{F}\leq Rn^{1+o(1)}$, which is almost the same as the other two cases.
@@
Lemma 5.3 means that, at a fixed timestamp $\tau$, if the weights $w(\tau)$ is close to its initialization, the loss is decreasing, and the gradient is also small, then we can conclude at timestamp $\tau+1$, these conditions still hold as local convexity proved by Lemma 5.1. Thus, after checking the initial condition, we can conclude Theorem 4.2.
5.1 Technical Novelty and Comparison to the Existing Literature
In this work, as we consider the softmax activation function, the denominator term will also contribute to gradient calculation. Handling the denominator poses many technical challenges, where these challenges are unique to our setting and not presented in previous settings as ReLU [135], or $\exp$ [48] activation function. In detail, in the gradient calculation, we need new loss decomposition Lemma E.1 to split the loss into $\|F(\tau+1)-Y\|_{F}^{2}=\|F(t)-Y\|_{F}^{2}+C_{0}+C_{1}+C_{2}+C_{3}$. Then, we need to bound these new terms in Lemma E.3 for $C_{0}$, Lemma E.4 and Claim E.5 for $C_{1}$, Claim E.6 for $C_{2}$ and Claim E.7 for $C_{3}$, where all these Lemmas are novel and non-trivial. We refer readers to Appendix C.3 for more details.
6 Extension on Diffusion
@@ 6.1 Preliminary of Diffusion
$x(t)=$ $~e^{-\int_{0}^{t}\frac{1}{2}g(s)\mathrm{d}s}x(0)$
$~+e^{-\int_{0}^{t}\frac{1}{2}g(s)\mathrm{d}s}\int_{0}^{t}e^{\int_{0}^{s}\frac{1}{2}g(u)\mathrm{d}u}\sqrt{g(s)}\mathrm{d}B_{s}.$
Backward Process. We denote y(t)=x(Tt)y(t)=x(T-t) to reverse the forward process in time [62, 47, 21] that transforms noise into samples from the target distribution. We have a backward process associated to Eq. (3) as:
$\mathrm{d}y(t)=$ $~(\frac{1}{2}g(T-t)y(t)+g(T-t)\nabla\log p_{T-t}(y(t)))\mathrm{d}$
$~+\sqrt{g(T-t)}\mathrm{d}\bar{B}_{t},\quad y(0)\sim q_{0}.$ (4)
where $(\bar{B}_{t})_{t\geq 0}$ is another $d$-dim Brownian motion/noise. Following the literature, we call $\nabla\log p_{t}(\cdot)$ as “score function” [133]. We have $q_{0}$ as the initial distribution of the backward process and the score function $\nabla\log p_{t}(\cdot)$ as the gradient of log density of $x(t)$.
However, In practice, Eq.(6.1) cannot be directly used as both the score function and the distribution $p_{T}$ are unknown. To solve the problem, we (1) randomly select a noise distribution as the initial distribution of the backward process $p_{T}$; (2) replace the ground-truth score function $\nabla\log p_{t}(x(t))$ by an estimator $s_{\theta}(x(t),t)$. The parameterized estimator $s_{\theta}$ is learned by a neural network such as U-Net [63, 115] and Transformer [107]. Thus, we obtain a practically implementable approximation of the backward SDE:
$\mathrm{~d}y(t)=$ $~(\frac{1}{2}g(T-t)y(t)+g(T-t)s_{\theta}(y(t),t))\mathrm{d}t$
$~+\sqrt{g(T-t)}\mathrm{d}\bar{B}_{t},\quad y(0)\sim\mathcal{N}(0,I_{d}),$
@@
Score Matching. When estimating the score function, we usually use $L_{2}$ loss between the estimated and actual score:
@@
$~\min_{\theta}\frac{1}{T-T_{0}}\cdot\int_{T_{0}}^{T}\lambda(t)$
$~\mathbb{E}[\|s_{\theta}(x(t),t)-\nabla\log p_{t\mid 0}(x(t)\mid x(0))\|_{2}^{2}]\mathrm{d}t.$ (6)
In practice, the estimator of the score function is parameterized by a neural network, and we have the following sampling procedure for any $i\in[n]$,
@@
and we get the training dataset $\{x(0)_{i},(t_{i},x(t_{i})_{i})\}_{i=1}^{n}$, where $x(0)_{i}\in\mathbb{R}^{d}$ and $(t_{i},x(t_{i})_{i})\in\mathbb{R}^{d+1}$. We denote $x(0)$ as the noisy label and $\operatorname*{{\mathbb{E}}}[x(0)|x(t)]$ as the true label. For simplicity, we denote $x(0)_{i}$ as $y_{i}\in\mathbb{R}^{d}$ and $(t_{i},x(t_{i})_{i})$ as $x_{i}\in\mathbb{R}^{d+1}$ and the training dataset as ${\cal D}_{n}=\{(x_{i},y_{i})\}_{i=1}^{n}$. Here, $y$ denotes the image from a dataset, and $x$ denotes the noised image with its diffusion process time $t$.
Neural Network Parameterization. Recall that we consider a two-layer network with softmax activation function as the diffusion model in Eq. (1), satisfying $\forall\ell\in[d]$, $F(W,x,a)_{\ell}={{}{}{}m}\langle a_{\ell},\exp(W^{\top}x)\rangle\cdot\langle\exp(W^{\top}x),{\bf 1}_{m}\rangle^{-1}$. Note that we do not train the top-layer weights $a$, so we can denote it as $F_{nn}(W,x)$.
Then, similar as [63, 59], our loss function Eq. (6.1) can be rewrite as
@@ 6.2 Main Result of Diffusion
The function $g$ is almost everywhere continuous and bounded on $0,\infty)$.
@@
$\Gamma_{\delta}:=$ $~\left(\frac{2d^{2}A(R_{\mathcal{H}})}{\lambda}\log^{3/2}(\frac{e(dn)^{3/2}A(R_{\mathcal{H}})}{\lambda})+\frac{1}{\sqrt{n}}\right)^{2}$
$~+\frac{d^{2}A^{2}(R_{\mathcal{H}})}{\lambda^{2}}(\log(1/\delta)+\log(\log n)).$
Assumption 6.5 (Assumption 3.11 in [59]).
Fix any $F_{\mathcal{H}}\in\mathcal{H}$ with $\|F_{\mathcal{H}}\|_{\mathcal{H}}^{2}\leq R_{\mathcal{H}}$ and assume labels are generated as $\widetilde{y}_{j}=F_{\mathcal{H}}(x_{j})+\epsilon_{j}$. Suppose $)\widetilde{F}_{ntk}(\gamma(\widehat{T}$$\cdot)$ is obtained by GD-trained kernel regression with the number of iterations $\widehat{T}$. We assume there exists $\epsilon$ such that
$~\frac{1}{T}\int_{0}^{T}\mathbb{E}[\widetilde{F}_{ntk}(\gamma(\widehat{T}),(t,x(t)))-F_{\mathcal{H}}(t,x(t))\|_{2}^{2}]\mathrm{d}t$
[…]
and $\epsilon(n,\widehat{T})\rightarrow 0$ as $n\rightarrow\infty.$
@@
[…]
$~\frac{1}{T}\int_{0}^{T}\lambda(t)\mathbb{E}[\|s_{W(\widehat{T})}(t,x(t))-\nabla\log p_{t}(X_{t})\|_{2}^{2}]\mathrm{d}t$
$\leq$ $~O(\frac{1}{\lambda\sqrt{n}}+\epsilon(n,\widehat{T})+dA^{2}(R_{\mathcal{H}})$
$~+dA(R_{\mathcal{H}})+\sqrt{dA(R_{\mathcal{H}})\Gamma_{\delta}}+\Gamma_{\delta}).$
Please refer to Appendix G.1 for the complete proof. Here, we provide a proof sketch.
@@ 7 Conclusion
This paper provides a theoretical analysis of the optimization and generalization properties of two-layer neural networks with the softmax activation function. We apply our results in learning score estimation functions in diffusion models with noisy labels to verify our analysis effectiveness. Our findings contribute to a deeper understanding of the power of softmax neural networks and their potential for self-attention, advanced LLMs, and generative modeling.
@@ Acknowledgement
Roadmap. In Section A, we introduce some definitions that will be used in the proof. In Section B, we provide the basic concentration. In Section C, we provide the proof of our inductions. In Section D, we establish a bound for the weight of induction Part 1. In Section E, we establish a bound for the loss of induction Part 2. In Section F, we introduce the NTK regression. In Section G, we introduce the diffusion. In Section H, we discuss the potential implications of our results for popular frameworks such as attention mechanisms and feature learning.
@@ C.1 Main Result
Then, after $\widehat{T}$ iterations, with probability at least $1-\delta$, we have
@@ E.3 Bounding $C_{0}$
$=$ $~{{}{}{}m}(\sum_{r=1}^{m}a_{\ell,r}\alpha_{i}(\tau+1)^{-1}\alpha_{i}(\tau)^{-1}\sum_{r_{2}=1}^{m}-\eta\langle\Delta w_{r_{2}}(\tau),x_{i}\rangle\exp(w_{r_{2}}(\tau)x_{i})\exp(\langle w_{r}(\tau+1),x_{i}\rangle)+\eta^{2}\Delta_{2})$
@@ Appendix F NTK Regression
Recall that, by Definition 3.6, we have the $H^{*}=H(W(0))$. Recall $H^{*}]_{i,j}\in\mathbb{R}^{d\times d}$ is the kernel between $x_{i}$ and $x_{j}$. Similarly, $\forall\ell_{1},\ell_{2}\in[d]$, for test data, we can define the NTK induced feature map as
@@ F.1 Equivalence between Trained Net and Kernel Regression
where we have used the fact that the initial prediction is 0 as $\beta(0)={\bf 0}_{d}$. Similarly for $F_{nn}(x_{te})_{\ell}$. Let $F_{ntk2,i}(\tau)=F_{ntk2}(\beta(\tau),x_{i})$ and $F_{ntk2}(\tau)\in\mathbb{R}^{d\times n}$. Similarly, for the NN predictor $F_{nn}$. Now we take a closer look at the time derivative:
@@
Since $H^{*}$ is positive semidefinite, $H^{*}\operatorname{vec}(F_{nn}(\tau)-F_{ntk2}(\tau))$ term only makes $\|F_{nn}(\tau)-F_{ntk2}(\tau)\|_{F}$ smaller. Therefore, we have
@@ Appendix G Diffusion
$Z_{4}=$ $\frac{1}{T}\int_{0}^{T}\mathbb{E}[\|F_{\mathcal{H}}(t,x(t))-F_{*}(t,x(t))\|_{2}^{2}]\mathrm{d}t.$ $\text{(approximation)}.$
@@ G.1 Main Result of Diffusion
[…]
@@
Thus, we finish the proof by Claim G.1, where coupling is from above, label mismatch is from Theorem G.4, early stopping is from Assumption 6.5 and approximation is from Theorem G.3. ∎
@@ G.2 Tools From Previous Works
We have the following statements from previous works [59].
@@
$~\frac{1}{T}\int_{0}^{T}\mathbb{E}[\|F_{ntk}(\gamma(\tau),(t,x(t)))-\widetilde{F}_{ntk}(\gamma(\tau),(t,x(t)))\|_{2}^{2}]\mathrm{d}t$
$~\leq dA(R_{\mathcal{H}})+C_{0}(\sqrt{dA(R_{\mathcal{H}})\Gamma_{\delta}}+\Gamma_{\delta})$
@@
Appendix H Discussion
In this section, we provide discussions about the potential extensions of our method on various popular frameworks, such as attention mechanism (Section H.1) and feature learning (Section H.2).
H.1 Self-attention Learning
The self-attention can be written as
$F(W^{K}X,W^{Q}X,W^{V}X)\in\mathbb{R}^{d\times n^{\prime}},$ (27)
H.2 Feature Learning
Recently, there is a line of work showing that feature learning may be beyond NTK on sample complexity or time complexity, e.g., [5, 148, 60, 2, 38, 23, 153, 70, 85, 56, 114, 92, 90, 37, 124, 125] and many more. It is worth studying the feature learning ability of two-layer softmax NN to figure out what feature pattern the softmax prefers to learn and how it happens. We leave this valuable direction as a future work.